OpenStudy (anonymous):

simplify the ratio 20 in. : 5 ft

Directrix (directrix):

20/ 60 = 1/3

Directrix (directrix):

1 foot = 12" so the ratio is 20 inches over (5 times 12) = 20/60 = 1/3

OpenStudy (anonymous):

i get it thanks

Directrix (directrix):

Glad to help

OpenStudy (anonymous):

:)